zarkam21:

Compare solids, liquids, and gases in terms of their shape when placed into different containers Liquids- The intermolecular attractive forces are strong enough to hold molecules close together Liquids are more dense and less compressible than gasses, Liquids have a definite volume, independent of the size and shape of their container solid - The intermolecular forces between neighboring molecules are strong enough to keep them locked in position, Solids are not very compressible due to the lack of space between molecules gas-widely separated molecules. The kinetic energy of the molecules is greater than any attractive forces between the molecules The lack of any significant attractive force between molecules allows a gas to expand to fill its container

Vocaloid:

Good, just mention that liquids do take the shape of their container

zarkam21:

A sample of gas is in a rigid container that cannot change volume. Use the kinetic molecular theory to explain why the pressure in the container will increase as the temperature increases.

zarkam21:

not sure about this

Vocaloid:

the pressure of a gas comes from the force exerted from the collisions of molecules against the walls of the container. higher temp --> higher molecular velocity --> higher collision frequency --> higher pressure

zarkam21:

Write one or two sentences to describe how heat will flow between two objects of different temperatures that are touching.

zarkam21:

If object A is hotter than object B, that means that the atoms have more energy. They are moving, colliding or simply vibrating. When object A comes in contact with object B, it will transfer the energy and set the neighboring particles also vibrating. Thus making it hot.

Vocaloid:

you're on the right track, just also mention that the heat is transferred until both objects are at the same temperature

zarkam21:

If there are two objects, object A and object b. And for example, if object A is hotter and higher in temperature than object b, it will result in more energy. These atoms are moving and colliding with each other. The heat is transferred until both objects are at the same temperature

Vocaloid:

good I'm going to make a quick commute to a new study space, I'll be back asap

zarkam21:

ok

zarkam21:

Write one or two sentences to compare the internal energy of a molecular gas with the internal energy of a monatomic gas.

zarkam21:

For a monatomic ideal gas, each of the three directions (x, y, and z) contribute 1/2 kT per molecule, for a total of 3/2 kT per molecule... and a polyatomic gas has contributions from three translation direction and three axes of rotation, giving U = 6/2 NkT, or 3NkT." So the polyatomic gas has more energy than a monatomic gas at the same temperature.

Vocaloid:

well done

zarkam21:

Which states of matter include the potential energy from intermolecular forces when measuring their internal energies?

zarkam21:

SOlid and liquid

Vocaloid:

good (I'm not sure about plasma, but I think solid + liquid would be an acceptable answer)

zarkam21:

What is the average velocity of atoms in 1.00 mol of neon (a monatomic gas) at 465 K? For m, use 0.0202 kg.

zarkam21:

847 m/s

Vocaloid:

nvm

zarkam21:

What is the internal energy of 3.00 mol of diatomic oxygen gas, O2, at 288 K

Vocaloid:

diatomic would be (5/2)n k T I believe

zarkam21:

Ugh stuck on this one

Vocaloid:

(5/2)nkT = (5/2)(3)(1.38064852 × 10^-23)(288) = ?

zarkam21:

umm 454

zarkam21:

i used mathway

Vocaloid:

oh whoops it's (5/2)nRT not kt (5/2)(3.00)(8.31)(288) = ?

zarkam21:

17949.6

Vocaloid:

good

zarkam21:

Water has many unique properties. Identify a property of water and explain why this property makes water vital to living organisms.

zarkam21:

high specific heat because

zarkam21:

of the high heatall types of animals can use it

zarkam21:

wait no how about attraction to polar molecules because

Vocaloid:

remember yesterday we said that water has a high specific heat which allows it to stay liquid over a high range of temperatures? can you think of why that might be important to animals?

zarkam21:

This is important because even if it is cold outside animals will be able to drink good temperature water which is important for them to live

Vocaloid:

good

zarkam21:

Write one to two sentences that summarize Bernoulli's principle.

zarkam21:

Bernoulli studied fluids in pipes and found that where the speed of a fluid increases, the internal pressure in the fluid decreases. the pressure in these narrower regions decreases within the fluid as the water speeds up. Likewise, the pressure in the fluid increases as the water slows down in the wider regions.

Vocaloid:

good

zarkam21:

. A force of 8800 N is exerted on a piston that has an area of 0.010 m^2. What is the force exerted by a second piston that has an area of 0.040 m^2?

Vocaloid:

pascal's equation F1/A1 = F2/A2 find F2

zarkam21:

8800/0.010m^2=F/0.040 m^2

Vocaloid:

good, keep going

zarkam21:

140800

Vocaloid:

check your calculations again just try doing 8800/0.01 = F/0.04

zarkam21:

35200

Vocaloid:

good

zarkam21:

An object's weight is equal to its buoyant force in water. What would you expect to happen when the object is placed in water? Explain your reasoning.

zarkam21:

the object will sink if the force than the floating pressure

zarkam21:

less*

Vocaloid:

there are two forces that determine whether an object floats/sinks, buoyant force and weight force, if weight force = buoyant force, the object is neutrally buoyant which means it neither sinks nor rises

Vocaloid:

anyway I need to start heading home, get back around 5:30 or so

zarkam21:

Sounds good